OpenStudy (anonymous):

Determine whether the following series is divergent, conditionally convergent or absolutely convergent.

OpenStudy (anonymous):

\[\sum_{n=1}^{∞}\frac{ n+(-1)^n7 }{ 2n^2+1 }\]

OpenStudy (anonymous):

according to my calculations, it is not absolutely convergent.

OpenStudy (anonymous):

So I need to show the series are convergent or divergent.

OpenStudy (amistre64):

did you try a comparison? for large n, this is just 1/2n

OpenStudy (amistre64):

err, |7/2n^2| maybe

OpenStudy (anonymous):

But the comparison are only for positive series, or what?

OpenStudy (amistre64):

|an| is a positive series .... but im just thinking

OpenStudy (amistre64):

do you know what a limit test is by chance?

OpenStudy (anonymous):

Yes but only if I have to show it is absolutely convergent.

OpenStudy (anonymous):

Yes I know the limit test..

OpenStudy (amistre64):

"Fractions involving only polynomials or polynomials under radicals will behave in the same way as the largest power of n will behave", lamars site ...

OpenStudy (amistre64):

does this sound reasonable for a comparison? \[\frac{(-7)^n}{2n^2}\]

OpenStudy (amistre64):

\[\frac{7}{2n^2}\frac{2n^2+1}{7+n}\] \[\frac{7}{7+n}\frac{2n^2+1}{2n^2}\] \[\frac{7+n-n}{7+n}\frac{2n^2+1}{2n^2}\] \[lim~(1-\frac{n}{n+7})(1+\frac{1}{2n^2})\] (1-1)(1+0) = 0 since the limit is not positive .... it diverges

OpenStudy (amistre64):

or at least thats the way im seeing it ....

OpenStudy (anonymous):

Okay. Sp there you show that the series are not absolutely convergent..

OpenStudy (amistre64):

i cant really be sure that this is correct, but i think that shows that 7/2n^2 is a viable comparison; and since that diverges .... is it smaller than the original one?

OpenStudy (amistre64):

ugh ... i do loathe these things

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

just forget everything ive posted so far ... i was trying to recall a few things that im sure i messed up

OpenStudy (anonymous):

Thank you anyway.